How do you find the answer for adding faction like 3/4+4/9
Answer by richard1234(7193)

Best to rewrite fractions using a common denominator.

LCM of denominators = LCM(4,9) = 36

3/4 = 27/36
4/9 = 16/36

3/4 + 4/9 = 27/36 + 16/36 = 43/36, which can also be written 1 7/36.